OpenStudy (anonymous):

Which pair of ratios are equivalent? 3/5 and 5/8 7/5 and 4/3 2/1 and 8/4 5/4 and 6/5

OpenStudy (king):

2/1 and 8/4

OpenStudy (ash2326):

2/1 and 8/4

OpenStudy (anonymous):

2/1 and 8/4

OpenStudy (anonymous):

2*4/1*4 = 8/4 so they are equivalent.

OpenStudy (anonymous):

2/1 and 8/4 8/4 =2 2/1 = 2